Mzee Kobe Bank wished to establish the times in seconds that each ATM transaction takes. A sample of ATM users were observed and the time in seconds each spent at the ATM was as follows:



Time  (seconds)



10-19



20-29



30-39



40-49



50-59



Number of customer



15



60



67



98



2



 



i. Calculate the coefficent of variation of the waiting time?  ( 6 marks)





1
Expert's answer
2021-11-02T12:40:36-0400

coefficent of variation:

"CV=\sigma/\mu"


"\mu=\sum(m_if_i)/N"

where m is midpoint of the class,

f is frequency.


"\mu=\frac{14.5\cdot15+24.5\cdot60+34.5\cdot67+44.5\cdot98+54.5\cdot2}{15+60+67+98+2}=35"


"\sigma=\sqrt{\frac{f_i(m_i-\mu)^2}{N}}=\sqrt{\frac{ 15(14.5-35)^2+60(24.5-35)^2+67(34.5-35)^2+98(44.5-35)^2+2(54.5-35)^2}{242}}=9.65"